OpenStudy (anonymous):

The angles of a triangle all add up to 180 degrees. If in a triangle, the measure of the largest angle 12 more than 5 times the measure of the smallest angle. The measure of the third angle is twice the measure of the smallest angle. What is the measure of the largest angle?

OpenStudy (anonymous):

I got 21 for the smallest and 117 for the largest? is this correct?

OpenStudy (phi):

if we use L (for largest), S (for smallest, and T (for third angle) then they say: largest angle is 12 more than 5 times the measure of the smallest angle. L is 12 more than 5*S --> L = 12 + 5 S third angle is twice the smallest angle T = 2 * S so we have L = 12 + 5 S T= 2S the sum of the 3 angles is L+S+T= 180 using our equations to replace L and T we write 12+5S + S + 2S = 180 or 8S +12 = 180 add -12 to both sides 8S = 168 divide both sides by 8 S= 21 and L = 12 + 5*21 = 12+105 = 117